A 1.000-L vessel contains a mixture of hydrogen, carbon monoxide and methane gases in unknown proportions at 23.5-bar and 25 °C. The gaseous mixture is found to have a mass of 12.70-g and a higher heating value of 353.40 kJ/mol. What is the composition of the gas?

y(H2) = ?

y(CO) = ?

y(CH4) = ?
Answer #1

n = PV/RT = (23.197x1)/(0.0821x298) = 0.948

nTotal = n H2 + nCO + n CH4   = 0.948

mH2 + mCO + mCH4 = 12.7 gram

(mH2)/2 + (mCO )/28 + (mCH4)/16 = 0.948 gram